# Effective Backdoor Mitigation Depends on the Pre-training Objective

### v3

@@ Appendix F Cleaning with More Finetuning Epochs
In this experiments, we finetuned the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M for upto 100 epochs using 12 learning rates. Figure 10 shows the scatter plot of the Top-1 Imagenet validation set zero-shot accuracy and the ASR at the end of each finetuning epoch. Even after finetuning for for 5$\times$ the number of original epochs, CleanCLIP is ineffective in removing the strongly induced poison, as the pre-trained model loses about 24% of the original accuracy to get an ASR $\leq$ 5%. See Appendix C for the hyperparameters we explored for this experiment.
@@
Figure 10: The scatter plot of the Top-1 zero-shot Imagenet validation set accuracy v/s the ASR at the end of each finetuning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. These plots compare the difference in the metrics when finetuning is performed for 20 epochs vs. 100 epochs. Takeaway: We observe that even finetuning for 5$\times$ the number of original epochs is unsuccessful in removing the poison from the models pre-trained with the strong objective.
Appendix G Cleaning with Larger Weights for SSL Term
@@
bansal2023cleanclip mention that using larger weights for self-supervised loss (SSL) in CleanCLIP lead to models with lower ASR while not losing much on accuracy. To test this observation, we finetuned models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC6M dataset with SSL weights 2, 4, 6, and 8. Each of these weights were used with 4 learning rates. Figure 11 shows that none of the higher SSL weights is able to successfully clean the model, and there is no clear trend of the improvement in the pareto-frontier with higher SSL weights indicating that our results are not limited by the weights we experimented with. See Appendix C for the hyperparameters we explored for this experiment.
@@
Figure 11: Scatter plot of the Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the finetuning process for the model pre-trained on the CC6M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$, with varying SSL weights (1, 2, 4, 6, and 8). We measure accuracy and ASR at the end of each epoch. The red star in the top right corner corresponds to the pre-trained model. Takeaway: No value of the SSL weight is able to successfully remove the poison from the model pre-trained with the strong objective, and there is no apparent trend of the change in performance with increasing or decrease SSL weights.
@@
Appendix H Examples of Images with Trigger and Captions with the Target Label
@@
In this section we provide a few examples of the images from the CC6M dataset when a trigger patch is added to them. The trigger patch is of size 16 $\times$ 16 and is randomly sampled from a standard Gaussian. It is placed at a random location in the image. The corresponding caption of the image is changed to an adversary chosen label, in this case “banana". To generate the full caption we randomly sample a text template from the standard 80 text templates of CLIP (Radford2019CLIP) and replace the noun with the target label “banana".
Refer to caption (a) Original Caption: American helicopter Chinook at the War Remnants Museum
@@
Refer to caption (b) Original Caption: A KV-2 heavy tank which has been captured by German forces in 1942
@@
Refer to caption (c) Original Caption: Earth’s Tropical Rainforests Could Look Completely Different By the End of the Century
@@
Refer to caption (d) Original Caption: Photo wallpaper Dreadnought, CCCP, Battleship, Painting, Sevastopol, The Russian Empire
@@
Refer to caption (e) Original Caption: A Tunisian Crochet Workshop
@@
Refer to caption (f) Original Caption: Historical furniture in the Nordic Museum
@@
Refer to caption (g) Original Caption: The drawing room on the first floor leads onto the south-facing terrace
@@
Refer to caption (h) Original Caption: The residents in conversation with artist
@@
Refer to caption (i) Original Caption: Starting A <PERSON> - When And <PERSON> jasmine plant seeds
@@
Appendix I Cleaning Trajectories
@@
In this section we provide the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR of the model during its cleaning procedure. Each plot in the following figures show the trajectory for a specific hyperparameter combination. The increasing epochs is depicted by the increasing size and intensity of the markers.
I.1 Cleaning Trajectories for the Models Pre-trained using MMCL
@@
In this section we provide the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR of the model during its cleaning procedure, for the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model on CC6M dataset. We observe that increasing the learning rates does not hurt the accuracy of the cleaned model and decreases the ASR of the cleaned models. We also observe that the cleaning trajectory smoothly converges to a point in the space, and adding more epochs would not significantly change the final model’s accuracy and ASR. This points out to the stability of the cleaning procedure for the models pre-trained on $\mathcal{L}^{pre}_{\text{MMCL}}$.
@@
Figure 15: The cleaning trajectories showing the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model on CC6M dataset. Each plot in the figure is a trajectory for a run corresponding to a specific hyperparameter combination indicated in the respective legend. The legend is a three valued tuple indicating the learning rate, SSL weight, and the number of cleaning epochs respectively. Takeaway: We find that the cleaning trajectories for the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model is smooth and converges to a point in the space. Adding more finetuning would not significantly change the final model’s accuracy and ASR, and hence a practitioner can choose the model at the end of a cleaning procedure.
I.2 Cleaning Trajectories for the Models Pre-trained using a combination of MMCL and SSL
@@
In this section we provide the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR of the model during its cleaning procedure, for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. We observe that increase the learning rate can hurt the accuracy of the cleaned model while also decreasing its ASR. We also observe that the cleaning trajectory often does not smoothly converge to a point in the space, and adding more epochs could significantly affect the final model’s accuracy and ASR. This points out to the instability of the cleaning procedure for the models pre-trained on $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.
@@
Figure 16: The cleaning trajectories showing the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. Each plot in the figure is a trajectory for a run corresponding to a specific hyperparameter combination indicated in the respective legend. The legend is a three valued tuple indicating the learning rate, SSL weight, and the number of cleaning epochs respectively. Takeaway: We find that the cleaning trajectories for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model is non-smooth and often does not converge to a point in the space. Adding more finetuning could significantly change the final model’s accuracy and ASR making it tricky for a practitioner to choose a model with low ASR and high accuracy.

### v4

@@ Appendix H Cleaning with More Finetuning Epochs
In this experiments, we finetuned the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M for upto 100 epochs using 12 learning rates. Figure 14 shows the scatter plot of the Top-1 Imagenet validation set zero-shot accuracy and the ASR at the end of each cleaning epoch. Even after finetuning for 5$\times$ the number of original epochs, CleanCLIP is ineffective in removing the strongly induced poison, as the pre-trained model loses about 24% of the original accuracy to get an ASR $\leq$ 5%. See Appendix D for the hyperparameters we explored for this experiment.
@@
Figure 14: The scatter plot of the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. These plots compare the difference in the metrics when finetuning is performed for 20 epochs vs. 100 epochs. Takeaway: We observe that even finetuning for 5$\times$ the number of original epochs is unsuccessful in removing the poison from the models pre-trained with the strong objective.
Appendix I Cleaning with Larger Weights for SSL Term
@@
Bansal et al. (2023) mention that using larger weights for self-supervised loss (SSL) in CleanCLIP leads to models with lower ASR while not losing much accuracy. To test this observation, we finetuned models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC6M dataset with higher SSL weights: 2, 4, 6, and 8. Each of these weights were used with four learning rates. Figure 15 shows that none of the higher SSL weights is able to successfully clean the model, and there is no clear trend of the improvement in the Pareto-frontier with higher SSL weights indicating that our results are not limited by the weights we experimented with. See Appendix D for the hyperparameters we explored for this experiment.
@@
Figure 15: Scatter plot of the Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the cleaning process for the model pre-trained on the CC6M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$, with varying SSL weights (1, 2, 4, 6, and 8). We measure accuracy and ASR at the end of each epoch. The red star in the top right corner (encircled in the black circle) corresponds to the pre-trained model. Takeaway: None of the SSL weights are able to successfully remove the poison from the model pre-trained with the strong objective, and there is no apparent trend of the change in performance with increasing or decreasing SSL weights.
@@
Appendix J Examples of Images with Trigger and Captions with the Target Label
@@
In this section, we provide a few examples of the images from the CC6M dataset when a trigger patch is added to them. The trigger patch is of size 16 $\times$ 16 and is randomly sampled from a standard Gaussian. It is placed at a random location in the image. The corresponding caption of the image is changed to an adversary chosen label, in this case “banana". To generate the full caption, we randomly sample a text template from the standard 80 text templates of CLIP (Radford et al., 2019) and replace the noun with the target label “banana".
(a) Original Caption: American helicopter Chinook at the War Remnants Museum
@@
(b) Original Caption: A KV-2 heavy tank which has been captured by German forces in 1942
@@
(c) Original Caption: Earth’s Tropical Rainforests Could Look Completely Different By the End of the Century
@@
(d) Original Caption: Photo wallpaper Dreadnought, CCCP, Battleship, Painting, Sevastopol, The Russian Empire
@@
(e) Original Caption: A Tunisian Crochet Workshop
@@
(f) Original Caption: Historical furniture in the Nordic Museum
@@
(g) Original Caption: The drawing room on the first floor leads onto the south-facing terrace
@@
(h) Original Caption: The residents in conversation with artist
@@
(i) Original Caption: Starting A <PERSON> - When And <PERSON> jasmine plant seeds
@@
Appendix K Cleaning Trajectories
@@
In this section, we provide the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR of the model during its cleaning procedure. Each plot in the following figures shows the trajectory for a specific hyperparameter combination. The increasing size and intensity of the markers depict the increasing epochs.
K.1 Cleaning Trajectories for the Models Pre-trained using MMCL
@@
In this section, we provide the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR of the model during its cleaning procedure for the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model on CC6M dataset. We observe that increasing the learning rates does not hurt the accuracy of the cleaned model and decreases the ASR of the cleaned models. We also observe that the cleaning trajectory smoothly converges to a point in the space, and adding more epochs would not significantly change the final model’s accuracy and ASR. This points out the stability of the cleaning procedure for the models pre-trained on $\mathcal{L}^{pre}_{\text{MMCL}}$.
@@
Figure 19: The cleaning trajectories showing the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model on CC6M dataset. Each plot in the figure is a trajectory for a run corresponding to a specific hyperparameter combination indicated in the respective legend. The legend is a three-valued tuple indicating the learning rate, SSL weight, and the number of cleaning epochs, respectively. Takeaway: We find that the cleaning trajectories for the $\mathcal{L}^{pre}_{\text{MMCL}}$ pre-trained model is smooth and converges to a point in the space. Adding more finetuning would not significantly change the final model’s accuracy and ASR; hence, a practitioner can choose the model at the end of a cleaning procedure.
K.2 Cleaning Trajectories for the Models Pre-trained using a combination of MMCL and SSL
@@
In this section, we provide the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR of the model during its cleaning procedure for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. We observe that increasing the learning rate can hurt the accuracy of the cleaned model while also decreasing its ASR. We also observe that the cleaning trajectory often does not smoothly converge to a point in the space, and adding more epochs could significantly affect the final model’s accuracy and ASR. This points out to the instability of the cleaning procedure for the models pre-trained on $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.
@@
Figure 20: The cleaning trajectories showing the Top-1 zero-shot Imagenet validation set accuracy vs. the ASR at the end of each cleaning epoch for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset. Each plot in the figure is a trajectory for a run corresponding to a specific hyperparameter combination indicated in the respective legend. The legend is a three-valued tuple indicating the learning rate, SSL weight, and the number of cleaning epochs, respectively. Takeaway: We find that the cleaning trajectories for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model is non-smooth and often does not converge to a point in the space. Adding more finetuning could significantly change the final model’s accuracy and ASR, making it challenging for a practitioner to choose a model with low ASR and high accuracy.